Cost Item Details
Food ingredients
$2 Per footlong
Labor cost per footlong
Labor $15.00/hour wage rate, each worker can make 12 sandwiches per hour
Credit card transaction fee
1% of sales
Electricity
$360 per month
Rent
Rent $1,200 per month
Franchise fee amortization Franchise and startup fees $36000 divided by 180 months
Royalty fee 8.0% Sales
Advertising fee 4.5% of Sales
Equipment leasing cost $600 per month

Assume that all transactions are paid for using a credit card and a $5 footlong deal is in place.

1. Questions Identify each of the listed costs of one footlong sub sandwich as either variable (per unit) or fixed.

2. What is the contribution margin of each $5 footlong sub sandwich?

3. With a $5 footlong deal, how many sandwiches should a Subway franchisee sell per month to break-even?

4. With a $5 footlong deal, how many sandwiches should a Subway franchisee sell per month to earn a profit of $2,000 per month?

5. If 3,000 sandwiches are sold in a certain month, what is the margin of safety (in dollar amount) for the month?

6. Assume the price of a footlong becomes $6. How many sandwiches should a Subway franchisee sell per month to break-even?
Answer #1

1.

Variable Fixed
Food Ingredients $ 2.00
Labor cost 1.25
Credit card transaction fee 0.05
Electricity $ 360
Rent 1,200
Franchise fee amortization 200
Royalty fee 0.40
Advertising fee 0.225
Equipment leasing cost 600
Totals $ 3.925 $ 2,360

2. Contribution margin per sandwich = $ ( 5.00 - 3.925 ) = $ 1.075.

3. Number of sandwiches to be sold to break-even = $ 2,360 / $ 1,075 = 2,195.35 units or 2,196 units

4. Number of sandwiches to be sold in order to earn a profit of $ 2,000 = $ ( 2,360 + 2,000 ) / $ 1.075 = 4,055.81 units

5. If 3,000 units are sold, margin of safety = ( 3,000 - 2,196 ) x $ 5 = $ 4,020

6. Break-even number of units = $ 2,360 / $ ( 6 - 3.925 ) = 1,137.35 units.